# On $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ and Frege Systems

### v5

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. Since there exists oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mysterious behind it and showing that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there are language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which is in the area of proof complexity. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ 4. Lazy-diagonalization against All ${\rm co}\mathcal{NP}$-machines
5. Showing That $L_{d}\in\mathcal{NP}$
@@ A Frege proof system $F$ is an inference system for propositional logic based on
We can give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of $L_{d}\in\mathcal{NP}$ in Section 5, $A$ is the following:
@@
and one of the main deductions of Theorem 5.2, besides showing that $A$ holds, is to show that $A$ implies $B$.
@@
Now we introduce the notion of $p$-simulation between two proof systems.
@@ 8.1 Lower Bounds for Frege Systems
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g. [CR79]). In fact, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g. [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems, until they are established for all abstract proof systems. However, our approach is the opposite, that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
@@ 9 Conclusions
thus settling a very important and long-standing conjecture in computational complexity theory. As a by-product, the important result shown in [Lin21] that $\mathcal{P}\neq\mathcal{NP}$ and $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ also follows. We also obtain other interesting consequences such as Corollary 6.
@@
One of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$ is basically the same as the one in the author’s recent work [Lin21]. In other words, we first enumerate all polynomial-time ${\rm co}\mathcal{NP}$-machines, then by lazy-diagonalization we show that there exists a language $L_{d}$ accepted by a universal nondeterministic Turing machine running within time $O(n^{k})$ for all $k\in\mathbb{N}_{1}$, but by no ${\rm co}\mathcal{NP}$-machines. The next is to use essentially the same technique developed in author’s work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine is in fact in $\mathcal{NP}$, thus establishing the main theorem of this paper.
@@
We have broken the so-called Relativization Barrier by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
@@
Note again that by Corollary 3 we in fact have shown $\mathcal{P}\neq{\rm co}\mathcal{NP}$ because if $\mathcal{P}={\rm co}\mathcal{NP}$ then $\mathcal{P}={\rm co}\mathcal{P}=\mathcal{NP}$, a contradiction to Corollary 3. As a matter of fact, we have also shown by Theorem 5 that there is ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete, i.e., the complexity class ${\rm co}\mathcal{NP}$ also has a rich structure.
@@
As mentioned in the introduction Section, the question of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomial bounded, thus answering an open problem in [Pud08]. As pointed out in Remark 8.2, we did the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 7 is that nondeterministic Turing machines can simulate Frege proofs. But, we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs. For more detailed information about Extended Frege Systems, consult the reference [CR79, CN10, Kra95].
@@
Lastly, there are many other interesting open questions in the field of proof complexity that we did not touch on in this paper. The interested reader is referred to [Bus12, CR74, CR79, CN10, Kra95, Kra19, Pud08, Raz03, Raz15].

### v6

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mysterious behind it and showing that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable, thus showing that the technique of lazy-diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there are language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which belongs to the area of proof complexity. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ 4. Lazy-diagonalization against All ${\rm co}\mathcal{NP}$-machines
5. Proving That $L_{d}\in\mathcal{NP}$
@@ A Frege proof system $F$ is an inference system for propositional logic based on
We can give a more intuitive explanation about the inference rule of Modus Ponens: For example, in the proof of Theorem 5.2 in Section 5 where we prove that $L_{d}\in\mathcal{NP}$, $A$ is the following:
@@
and one of the main deductions of Theorem 5.2, besides showing that $A$ is true, is to show that $A$ implies $B$.
@@
Now, we introduce the notion of $p$-simulation between two proof systems.
@@ 8.1 Lower Bounds for Frege Systems
Initially, propositional proof complexity has been primarily concerned with proving lower bounds (even conditional) for the length of proofs in propositional proof systems which is extremely interesting and well justified in its own right, with the ultimate goal of settling whether $\mathcal{NP}={\rm co}\mathcal{NP}$ (see e.g. [CR79]). In fact, as mentioned in Section 1, such research directions are called Cook’s Program for separating $\mathcal{NP}$ and ${\rm co}\mathcal{NP}$ (see e.g. [Bus12]): Prove superpolynomial lower bounds for proof lengths in stronger and stronger propositional proof systems, until they are established for all abstract proof systems. However, our approach is “to do the opposite”, that is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems.
@@ 9 Conclusions
thus resolving a very important and long-standing conjecture in computational complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$-machines is enumerable. Then, we constructed a universal nondeterministic Turing machine to simulate all ${\rm co}\mathcal{NP}$-machines and to do the lazy-diagonalization operation to produce a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in author’s recent work [Lin21] to show that the language $L_{d}$ is in fact in $\mathcal{NP}$, thus establishing the desired conclusion of this paper.
@@
We stress that one of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$, is basically the same as the one developed in the author’s recent work [Lin21].
We have broken the so-called Relativization Barrier. Since there exists some oracle $A$ such that
$\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A},$
@@
which suggests that ordinary diagonalization techniques (lazy-diagonalization) are not capable of proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, see e.g. [BGS75]. We explored the mysterious behind this by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}$-machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques (i.e., lazy-diagonalization) will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
As a by-product, the important result that $\mathcal{P}\neq\mathcal{NP}$ which is shown in [Lin21], also follows. Furthermore, the conclusion of $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ follows from the similar proof of that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$.
@@
Note again that by Corollary 3 we in fact have shown $\mathcal{P}\neq{\rm co}\mathcal{NP}$ because if $\mathcal{P}={\rm co}\mathcal{NP}$ then $\mathcal{P}={\rm co}\mathcal{P}=\mathcal{NP}$, a contradiction to Corollary 3. As a matter of fact, we have also shown by Theorem 5 that there exists ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete, i.e., the complexity class ${\rm co}\mathcal{NP}$ also has a rich structure. We have also obtained other interesting consequences such as Corollary 6 which states that there is no super proof system.
@@
As mentioned in the introduction Section, the problem of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomial bounded (Theorem 7), thus answering an open problem in [Pud08]. As pointed out in Remark 8.2, we do the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 7 is that nondeterministic Turing machines can simulate Frege proofs. But, we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs. For more details about Extended Frege Systems, please consult the reference [CR79, CN10, Kra95].
@@
Lastly, there are many other interesting open problems in the field of proof complexity that we did not touch on in this paper, and the interested reader is referred to these references [Bus12, CR74, CR79, CN10, Kra95, Kra19, Pud08, Raz03, Raz15] for those excellent open problems.